For the reaction A+B+C→D+E, the initial reaction rate was measured for various initial concentrations of reactants. The following data were collected:

Trial [A]
(M)
[B]
(M)
[C]
(M)
Initial rate
(M/s)
1 0.40 0.40 0.40 1.2×10−4
2 0.40 0.40 1.20 3.6×10−4
3 0.80 0.40 0.40 4.8×10−4
4 0.80 0.80 0.40

4.8×10−4

Rate law equation

The rate of a chemical reaction depends on the concentrations of the reactants. For the general reaction between A and B,

aA+bB⇌cC+dD

The dependence of the reaction rate on the concentration of each reactant is given by the equation called the rate law:

rate=k[A]m[B]n

where k  is a proportionality constant called the rate constant. The exponent m  determines the reaction order with respect to A, and n determines the reaction order with respect to B. The overall reaction order equals the sum of the exponents (m+n).

Part A

What is the reaction order with respect to A?

Express your answer as an integer.

Part B

What is the reaction order with respect to B?

Express your answer as an integer.

Part C

What is the reaction order with respect to C?

Express your answer as an integer.
Answer #1

A + B + C DTE Rate law is given by 8 = K (A] [B] [co- respect to A, B & C. where r = rat eaction K= Rate constant m,no = OrdeComparing O and ☺ = 10.41 0.4 1.2 x 10 4.8 x 10-4 0.8 0.4 m=2. [A] Order with B) order with c) order with I rate law respect

Overall Overall = m+n+o =2+0+1 = 3